OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, −3) and (3, −5). −1 1 0 one half

OpenStudy (photon336):

are you familiar with the formula for slope?

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (alexandervonhumboldt2):

use slope formula: (y1-y2)/(x1-x2) or (y2-y1)/(x2-x1) you can use both of them.

OpenStudy (anonymous):

This is a practice question for me i just want to see how it works

OpenStudy (alexandervonhumboldt2):

substitute points (-3+5)/(1-3)=2/(-2)=-2/2=???

OpenStudy (photon336):